OpenStudy (anonymous):

Practicing using quadratic formula can someone check my work for errors -x^4+7= x^2-2

OpenStudy (anonymous):

-x^4-x^2+9=0 -1(x^4+x^2-9)=0\[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] \[x=\frac{ -1\pm \sqrt{1^{2}-4(1)(-9)} }{ 2(1) }\]

OpenStudy (anonymous):

\[x=\frac{ -1\pm \sqrt{1-36} }{ 2 }\]

OpenStudy (anonymous):

\[x=\frac{ -1\pm \sqrt{-35} }{ 2 }\]

OpenStudy (littlebird):

-4(1)(-9) two negatives make a positive

OpenStudy (xapproachesinfinity):

you were solving for x^2 so you need to find for x!

OpenStudy (xapproachesinfinity):

and he is right typo there

OpenStudy (littlebird):

You should have noticed something was wrong when you got -35 under the square root. You cannot find the square root of a negative number.

OpenStudy (anonymous):

yeah i see that now forgot to change my sign

OpenStudy (xapproachesinfinity):

no you can have negative suare root if you are dealing with complex system

OpenStudy (xapproachesinfinity):

you sill need to solve for x you have done x^2

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[x=\frac{ -1\pm \sqrt{1^{2}-4(1)(-9)} }{ 2(1) }\]

OpenStudy (anonymous):

\[x=\frac{ -1\pm \sqrt{1-4(-9)} }{ 2 }\]

OpenStudy (anonymous):

\[x=\pm \sqrt{\frac{ -1+37 }{ 2 }}\]

OpenStudy (littlebird):

?

OpenStudy (littlebird):

I'm getting confused.

OpenStudy (anonymous):

whered you get a -3 ffrom?

OpenStudy (littlebird):

I think I did something wrong in the beginning...meh

OpenStudy (littlebird):

Hopefully this fixes me and my crazy math

OpenStudy (littlebird):

I haven't gotten enough sleep lately. Math at 1 AM does not bode well for me...

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ -x^4+7= x^2-2}\) \(\normalsize\color{black}{ -x^4+7 \color{red}{-x^2+2}= x^2-2\color{red}{-x^2+2}}\) \(\normalsize\color{black}{ -x^4+7-x^2+2=0}\) \(\normalsize\color{black}{ -x^4-x^2+9=0}\) \(\normalsize\color{black}{ -1(x^4+x^2-9)=0}\) \(\normalsize\color{black}{ x^4+x^2-9=0}\) Nowe, let x² = s \(\normalsize\color{black}{ (x^2)^2+x^2-9=0}\) \(\normalsize\color{black}{ s^2+s-9=0}\) see what I am doing ?

OpenStudy (solomonzelman):

then you can use the quadratic formula :) Let me know if you need more help.

OpenStudy (anonymous):

sorry @SolomonZelman yes i was trying to avoid changing variables because im not eaxactly sure how to put em back so it =x and not s \[\frac{ -1 \pm \sqrt{1^{2}-4(1)(-9)} }{ 2(1)^2}\]

OpenStudy (anonymous):

\[\frac{ -1 \pm \sqrt{1+36} }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ -1\pm \sqrt{37} }{ 2 }\]

OpenStudy (anonymous):

and would that be a final answer?

OpenStudy (solomonzelman):

this is you soled for s, getting \(\LARGE\color{black}{ s= \frac{1±\sqrt{37}}{2} }\)

OpenStudy (solomonzelman):

This is just for the s

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ s= \frac{1+\sqrt{37}}{2} }\) and \(\LARGE\color{black}{ s= \frac{1-\sqrt{37}}{2} }\) but, s= x²

OpenStudy (solomonzelman):

So, \(\LARGE\color{black}{ x= (~\frac{1± \sqrt{37}}{2}~)^2 }\)

OpenStudy (anonymous):

would the 37 stay and sqrt canel and everything else is sqrd?

OpenStudy (anonymous):

\[x=\frac{ 2+37 }{ 4 }\] \[x=\frac{ 39 }{ 4 }\]

OpenStudy (anonymous):

or stay how you have it

OpenStudy (solomonzelman):

you can rewrite the answer if you want to. You were a bit off when you re-wrote it. ( 1+√37 )² / 2² ( 1+2√37+ 37 ) / 2² ( 2√37+ 38 ) / 4 ( √37+ 19 ) / 2 <▬▬ and ( 1-√37 )² / 2² ( 1-2√37 +37) / 2² ( -2√37 +38) / 4 (- √37 +19) / 2

OpenStudy (solomonzelman):

I would put the square roots at the end, although it doesn't matter at all. like this, (19 + √37) / 2 and (19 - √37) / 2 in other words, (19 ± √37) / 2

OpenStudy (anonymous):

im a bit confused right now why wouldnt the exponent cancel the sqrt?

OpenStudy (solomonzelman):

you are raising (1+√37) /2 t the second power, not just √37/2 to the second power.

OpenStudy (anonymous):

yes everything is being raised to the second power like (1+2)^2 what i dont understand is 1^2=1 37^2=1369 sqrt of 1369=37 2^2=4 wouldnt it be

OpenStudy (anonymous):

38/4?

OpenStudy (anonymous):

or do you not solve it

OpenStudy (solomonzelman):

you have 2 solutions for s, and they are \(\Large\color{blue}{\frac{1+\sqrt{37}}{2} }\) and \(\Large\color{blue}{\frac{1-\sqrt{37}}{2} }\) right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry this is confusing for me im getting better though :)

OpenStudy (solomonzelman):

So, and we know that s=x², because we substituted s instead of x². So to find x, we substitute each of the solutions for s, into s=x². Right ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\Large\color{blue}{x^2=\frac{1±\sqrt{37}}{2} }\) Sorry, I was wrong. I don't need to raise both sides to the second power, I take their square root.

OpenStudy (solomonzelman):

\(\Large\color{blue}{x^2=\frac{1±\sqrt{37}}{2} }\) \(\Large\color{blue}{x^2=\frac{1-\sqrt{37}}{2} }\) and \(\Large\color{blue}{x^2=\frac{1+\sqrt{37}}{2} }\) \(\Large\color{blue}{x=±\sqrt{ \frac{1-\sqrt{37}}{2} }}\) and \(\Large\color{blue}{x=± \sqrt{ \frac{1+\sqrt{37}}{2} }}\)

OpenStudy (anonymous):

ahh i forgot to sqrt it all

OpenStudy (anonymous):

no i did that just didnt put in the sqrt sign for 37

OpenStudy (solomonzelman):

Then you would say, \(\Large\color{blue}{x=\sqrt{ \frac{1-\sqrt{37}}{2} }}\) and \(\Large\color{blue}{x=\sqrt{ \frac{1+\sqrt{37}}{2} }}\) \(\LARGE\color{blue}{x= \frac{\sqrt{1-\color{white}{s}\sqrt{37}}}{\sqrt{2}} }\) and \(\LARGE\color{blue}{x=\frac{\sqrt{1-\color{white}{s}\sqrt{37}}}{\sqrt{2}} }\)

OpenStudy (solomonzelman):

the last thing, bottom right corner of my last reply, should be +√37

OpenStudy (solomonzelman):

Then multiply top and bottom times √2. In each solution.

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{x= \frac{\sqrt{2}(\sqrt{1-\color{white}{s}\sqrt{37}}~~)}{2} }\) and \(\LARGE\color{blue}{x=\frac{\sqrt{2} (\sqrt{1+\color{white}{s}\sqrt{37}}~~)}{2} }\)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{x= \frac{\sqrt{2-\color{white}{s}2\sqrt{37}}}{2} }\) and \(\LARGE\color{blue}{x= \frac{\sqrt{2+\color{white}{s}2\sqrt{37}}}{2} }\)

OpenStudy (solomonzelman):

And each fraction is ± .

OpenStudy (anonymous):

so \[x=\sqrt{74}\]\[x=\sqrt{4\sqrt{74}}\]

OpenStudy (anonymous):

or keep as fractions?

OpenStudy (solomonzelman):

you are not getting the correct answers. Where do you get √74 from ? 2√37 = √4 × √37 = √148 And this is not necessary. I posted the exact answer and how to obtain it.

OpenStudy (anonymous):

oh alright i thought i had to solve for each one im sorry and i got 74 canceling out 2's and multiplying 37 by 2

OpenStudy (solomonzelman):

yeah.... but nothing cancels though.

OpenStudy (xapproachesinfinity):

\(\Large\color{blue}{x=±\sqrt{ \frac{1-\sqrt{37}}{2} }}\) there should 4 possible answers

OpenStudy (xapproachesinfinity):

you have two that are + or minus don't forget that